# From What to Why: A Multi-Agent System for Evidence-based Chemical Reaction Condition Reasoning

### v1

@@ Abstract
The chemical reaction recommendation is to select proper reaction condition parameters for chemical reactions, which is pivotal to accelerating chemical science. With the rapid development of large language models (LLMs), there is growing interest in leveraging their reasoning and planning capabilities for reaction condition recommendation. Despite their success, existing methods rarely explain the rationale behind the recommended reaction conditions, limiting their utility in high-stakes scientific workflows. In this work, we propose ChemMAS, a multi-agent system that reframes condition prediction as an evidence-based reasoning task. ChemMAS decomposes the task into mechanistic grounding, multi-channel recall, constraint-aware agentic debate, and rationale aggregation. Each decision is backed by interpretable justifications grounded in chemical knowledge and retrieved precedents. Experiments show that ChemMAS achieves 20–35% gains over domain-specific baselines and outperforms general-purpose LLMs by 10–15% in Top-1 accuracy, while offering falsifiable, human-trustable rationales, which establishes a new paradigm for explainable AI in scientific discovery.
@@ 1 Introduction
With the rapid development of large language models (Naveed et al., 2025; Zhao et al., 2023) (LLMs), there has been a growing interest in leveraging their powerful reasoning and planning abilities for reaction condition recommendation (Bran et al., 2025). Current LLM-based approaches can be broadly categorized into retrieval-based (Zhang et al., 2024b; Chen et al., 2023) and reasoning-based approaches. Retrieval-based approaches search for similar reactions from external databases and transfer their conditions to the query reaction, which is usually enhanced by learned molecular embeddings or unsupervised chemical priors to improve retrieval quality (Andronov et al., 2023). In contrast, reasoning-based approaches directly prompt or fine-tune LLMs to infer suitable reaction conditions from molecular structures or textual descriptions (Qian et al., 2023; Zhou et al., 2025), and achieve improved zero-shot and few-shot generalization capabilities.
@@
By shifting from mere top-$k$ ranking to interpretable, evidence-backed reasoning, ChemMAS offers a new paradigm: one that is not only predictive but also justifiable, auditable, and suitable for closed-loop experimentation. In our evaluation, ChemMAS outperforms specialized chemical models (e.g., RCR (Gao et al., 2018), Reagent Transformer (Andronov et al., 2023)) by 20-30% Top-1 accuracy and surpasses leading general-purpose LLMs (e.g., GPT-5, Gemini 2.5) by 10-15% on average, validating its effectiveness and robustness.
@@
  • We benchmark ChemMAS against specialized chemical models and cutting-edge general-purpose LLMs, showing state-of-the-art performance with up to 30-point gains in Top-1 accuracy and robust generalization across diverse condition types.
@@ 2.4 Multi-Channel Recall
We maintain a structured Reaction Base $\mathcal{D}=\{(\tau_{n},\mathbf{r}_{n},\mathbf{p}_{n},\mathbf{c}_{n})\}_{n=1}^{N}$, where each entry contains the reaction type $\tau_{n}$, molecular representations of reactants $\mathbf{r}_{n}$ and products $\mathbf{p}_{n}$, and a condition triple $\mathbf{c}_{n}=(\mathrm{cat},\mathrm{sol},\mathrm{reag})$. Given the current reaction context $(\hat{\tau},\mathcal{R},\mathcal{P})$ from Memory, we perform three parallel queries including type-, reactant-, and product-centric, to obtain candidate index sets $\mathcal{S}_{t},\mathcal{S}_{r},\mathcal{S}_{p}$ (exact type match for $\mathcal{S}_{t}$, top-$k$ nearest neighbors by functional-group, MCS, and embedding similarity for $\mathcal{S}_{r}$ and $\mathcal{S}_{p}$). Without any scoring or rank fusion, an entry is admitted into Matched Conditions if it hits on any of the three tags. We define the unified retrieval result as the deduplicated union:
@@ 5.1 Main Results
As shown in Table 1, ChemMAS surpasses both specialized chemical models and state-of-the-art LLMs across all reaction types and Top-$k$ settings. It achieves relative Top-1 accuracy improvements ranging from 70% to over 90% when compared to domain-specific baselines such as RCR, Reagent Transformer, and MM RCR. Even against top-tier general-purpose LLMs like GPT-5 and Gemini 2.5-Pro, ChemMAS yields consistent relative gains of 15–25% in Top-1 accuracy, underscoring its strength in fine-grained mechanistic reasoning.
ChemMAS achieves strong Top-5 accuracies across all reaction types, ranging from 83.2% to 93.9%, and peaking on solvent1. Its particularly strong performance on challenging categories such as catalyst and solvent2 where many baselines struggle to surpass 65% Top-1 accuracy, further highlights the benefits of its multi-agent coordination and domain-aware tool integration. These results demonstrate that ChemMAS not only outperforms domain-specific systems, but also exceeds the capabilities of cutting-edge general LLMs.
Table 1: Top-k accuracy (%) across five types of reaction conditions: catalyst, solvent1, solvent2, reagent1, reagent2. We report accuracy at $k\in\{1,5,10\}$, corresponding to Top-1, Top-5, and Top-10 ranks. The proposed ChemMAS achieves the best performance across all settings.
Model Top-$k$ Accuracy (%)
Catalyst Solvent1 Solvent2 Reagent1 Reagent2
@@
Method Top-$k$ Accuracy (%)
Catalyst Solvent 1 Solvent 2 Reagent 1 Reagent 2
@@
Table 3: Ablation on the SFT and RL. The best and second-best results are bolded and underlined.
Top-$k$ Accuracy (%)
Training
Framework
Catalyst Solvent 1 Solvent 2 Reagent 1 Reagent 2
@@
Figure 4: Multi-agent ablation: Top-1 accuracy improvements across Catalyst, Solvent1/2, and Reagent1/2 when adding specialized agents on top of $\mathcal{A}_{Gen}$$\mathcal{A}_{Full}$.
@@ 5.2 Additional Quantitative Analysis
We conducted an ablation study to analyze the contribution of different components in ChemMAS. The ablation settings are as follows: (1) w/o Main FG, w/o By-Product, and w/o Reaction Type denote removing the corresponding elements from the Memory module; (2) w/o Multi-Agent Debate replaces multi-agent collaboration with a single-agent reasoning process, thereby eliminating conversational exchanges; (3) w/o Multi-Step Reasoning removes the iterative evidence-based reasoning chain within each agent, such that agents can only rely on prior knowledge and inter-agent debate without tool invocation; (4) w/o Candidate Pairing discards the pairwise elimination mechanism for candidate conditions, instead applying a global scoring and ranking procedure to directly select the top-50 candidates. As illustrated in Table 2, removing key components leads to substantial performance drops, underscoring their critical role in ChemMAS. Specifically, removing Main FG from the Memory module results in a significant decrease in performance, with an average drop of +8.4% across all reaction conditions, highlighting the crucial role of functional group extraction and analysis in reaction condition prediction. Similarly, removing Multi-Step Reasoning causes an average accuracy decrease of 12.3%, underscoring the importance of evidence-based multi-round reasoning.
To evaluate the effectiveness of each stage in our training framework, we conducted an ablation study focusing on SFT and RL. As illustrated in Table 3. The results demonstrate that removing either the SFT or RL stage leads to a clear degradation in Top-k Accuracy across all reaction condition types. Notably, excluding SFT results in a slightly larger drop compared to removing RL, underscoring the importance of supervised fine-tuning in establishing a strong initialization for subsequent reinforcement learning. These findings highlight the necessity of the two-stage training framework, where SFT and RL play complementary roles in achieving optimal performance.
@@ 6 Conclusion
We introduce ChemMAS, a multi-agent system that reframes reaction condition recommendation as the evidence-based reasoning task, grounded in domain-specific chemistry knowledge, mechanistic constraints, and interpretable evidence. Unlike prior models that focus solely on prediction, ChemMAS explains why each condition is appropriate, enhancing trust, generalization, and scientific utility. Empirically, ChemMAS achieves up to 30% improvements in Top-1 accuracy over specialized chemical models and consistently outperforms leading general-purpose LLMs. These results affirm the importance of transitioning from black-box predictions to evidence-based, auditable decision-making in scientific AI. In future work, we envision extending this agent-based reasoning framework to broader scientific domains such as materials design, bioinformatics, and physical simulation, where interpretability and mechanistic grounding are equally critical.
@@
    • Appendix C.1: Multi-Agent Debate Procedure
@@
  • Appendix D outlines the training and evaluation setups, including:
    • Appendix D.1: Optimization settings for ChemMAS and baselines
    • Appendix D.1: Agent-specific training targets and tool usage
    • Appendix D.2: Evaluation protocol and candidate selection details
    • Appendix D.3: Prompt design for different agent types
  • Appendix E presents additional experimental results and visualizations:
    • Appendix E.1: Top-5 and Top-10 Accuracy Ablation
    • Appendix E.2: Reaction prediction visualization table
@@ B.1 Reaction Condition Prediction
Predicting reaction conditions from reactants and products is a long-standing challenge in computer-aided synthesis. Early large-scale efforts such as (Gao et al., 2018) used feedforward neural networks trained on millions of Reaxys records to jointly predict catalysts, solvents, reagents, and temperatures, achieving promising top-k accuracies despite sparsity and label imbalance. Focusing on cross-coupling families, (Maser et al., 2021) formulated the task as multi-label ranking, developing role-specific encoders and leveraging graph-based features to yield accurate, context-aware predictions. To improve generalization and interpretability, (Wang et al., 2023) released benchmark datasets and proposed Parrot, a Transformer model augmented with unsupervised reaction center learning. Parrot achieved significant gains in condition accuracy and temperature estimation while offering interpretable attention maps localized to reactive substructures. Separately, (Andronov et al., 2023) addressed data quality limitations by training a Molecular Transformer to impute missing reagents in USPTO reactions. Their system not only improved reagent recall but also enhanced downstream product prediction models.
Retrieval-augmented methods incorporate external knowledge to improve robustness. TextReact (Qian et al., 2023) pairs structure-based encoders with retrieved literature snippets to inform condition prediction and retrosynthesis. By integrating textual context into training, it significantly outperforms molecule-only baselines. In peptide catalysis design, (Edwards et al., 2022) proposed a semi-automated ML framework for selecting universal catalyst libraries and discovered novel, high-selectivity peptides via efficient search in a large tripeptide space. At the interface of language and chemistry, (Edwards et al., 2022) introduced MolT5, a pre-trained encoder-decoder model that translates between molecules and natural language. It supports molecule-to-caption generation and chemically constrained text-to-molecule synthesis, offering a foundation for LLM-based explainability. More recently, (Zhang et al., 2024b) proposed a text-augmented multimodal LLM framework for reaction condition recommendation. Their method jointly encodes SMILES, molecular graphs, and relevant text to achieve state-of-the-art accuracy across open benchmarks and improve generalization under low-data or OOD settings. Despite these advances, current methods primarily focus on recommending what the potential reaction conditions are, but fail to provide explanatory why-level evidence for why such conditions are important or mechanistically justified.

### v2

@@ Abstract
The chemical reaction recommendation is to select proper reaction condition parameters for chemical reactions, which is pivotal to accelerating chemical science. With the rapid development of large language models (LLMs), there is growing interest in leveraging their reasoning and planning capabilities for reaction condition recommendation. Despite their success, existing methods rarely explain the rationale behind the recommended reaction conditions, limiting their utility in high-stakes scientific workflows. In this work, we propose ChemMAS, a multi-agent system that reframes condition prediction as an evidence-based reasoning task. ChemMAS decomposes the task into mechanistic grounding, multi-channel recall, constraint-aware agentic debate, and rationale aggregation. Each decision is backed by interpretable justifications grounded in chemical knowledge and retrieved precedents. Experiments show that ChemMAS achieves 20–35% gains over domain-specific baselines and outperforms general-purpose LLMs by 10–15% in Top-1 similarity, while offering falsifiable, human-trustable rationales, which establishes a new paradigm for explainable AI in scientific discovery.
@@ 1 Introduction
With the rapid development of large language models (LLMs) (Naveed et al., 2025; Zhao et al., 2023), there has been a growing interest in leveraging their powerful reasoning and planning abilities for reaction condition recommendation (Bran et al., 2025). Current LLM-based approaches can be broadly categorized into retrieval-based (Zhang et al., 2024b; Chen et al., 2023) and reasoning-based approaches. Retrieval-based approaches search for similar reactions from external databases and transfer their conditions to the query reaction, which is usually enhanced by learned molecular embeddings or unsupervised chemical priors to improve retrieval quality (Andronov et al., 2023). In contrast, reasoning-based approaches directly prompt or fine-tune LLMs to infer suitable reaction conditions from molecular structures or textual descriptions (Qian et al., 2023; Zhou et al., 2025), and achieve improved zero-shot and few-shot generalization capabilities.
@@
By shifting from mere top-$k$ ranking to interpretable, evidence-backed reasoning, ChemMAS offers a new paradigm: one that is not only predictive but also justifiable, auditable, and suitable for closed-loop experimentation. In our evaluation, ChemMAS outperforms specialized chemical models (e.g., RCR (Gao et al., 2018), Reagent Transformer (Andronov et al., 2023)) by 20-30% Top-1 similarity and surpasses leading general-purpose LLMs (e.g., GPT-5, Gemini 2.5) by 10-15% on average, validating its effectiveness and robustness.
@@
  • We benchmark ChemMAS against specialized chemical models and cutting-edge general-purpose LLMs, showing state-of-the-art performance with up to 30-point gains in Top-1 similarity and robust generalization across diverse condition types.
@@ 2.4 Multi-Channel Recall
We maintain a structured Reaction Base $\mathcal{D}=\{(\tau_{n},\mathbf{r}_{n},\mathbf{p}_{n},\mathbf{c}_{n})\}_{n=1}^{N}$, where each entry contains the reaction type $\tau_{n}$, molecular representations of reactants $\mathbf{r}_{n}$ and products $\mathbf{p}_{n}$, and a condition triple $\mathbf{c}_{n}=(\mathrm{cat},\mathrm{sol},\mathrm{reag})$. Given the current reaction context $(\hat{\tau},\mathcal{R},\mathcal{P})$ from Memory, we perform three parallel queries including type-, reactant-, and product-centric, to obtain candidate index sets $\mathcal{S}_{t},\mathcal{S}_{r},\mathcal{S}_{p}$ (exact type match for $\mathcal{S}_{t}$, top-$k$ nearest neighbors by functional-group, MCS, and embedding accuracy for $\mathcal{S}_{r}$ and $\mathcal{S}_{p}$). Without any scoring or rank fusion, an entry is admitted into Matched Conditions if it hits on any of the three tags. We define the unified retrieval result as the deduplicated union:
@@ 4.1 Training and evaluation setting
We measure performance using Top-$k$ Similarity, defined as the maximum Tanimoto similarity between the ground truth and top-$k$ predicted candidates, averaged over a composite of molecular fingerprints (Path-based, MACCS, Morgan). This metric reflects the best structural match retrieved by the model. Details are in the Appendix.
@@ 4.2 Datasets
Furthermore, we incorporated the RCR subset of ChemCoTBench (Li et al., 2025) as a lightweight public benchmark for small-scale evaluation. This subset of 90 high-quality, well-structured reaction–condition QA instances allows us to assess system generalization and stability under distribution shift. Further details on the private dataset and the ChemCoTBench-RCR subset are provided in the Appendix.
@@ 5.1 Main Results
Table 1: Generalization evaluation on ChemCoTBench. Top-$k$ similarity (%) for $k\in\{1,5,10\}$. The best and second-best results are bolded and underlined. Green values in parentheses show relative improvements over the second-best results.
Top-$k$ Similarity (%)
Model Catalyst Solvent Reagent
1 5 10 1 5 10 1 5 10
Zero-shot LLMs
Qwen3-235B-A22B 40.1 53.1 58.6 36.4 41.1 52.9 36.4 50.2 58.7
GPT5 41.9 59.2 66.1 44.1 57.5 65.2 40.1 55.1 61.1
Claude3.7-Sonnet 38.5 56.2 59.1 40.4 52.1 61.2 34.3 48.0 54.3
DeepSeek-R1 39.7 55.6 62.0 38.4 48.3 56.3 35.2 47.6 55.4
Gemini2.5-Pro 45.6 62.1 69.5 42.1 58.6 71.2 38.9 52.1 59.8
ChemMAS 62.1 68.3 76.1 57.8 66.5 76.8 51.2 59.1 67.7
(+16.5) (+6.2) (+6.6) (+13.7) (+7.9) (+5.6) (+11.1) (+4.0) (+6.6)
Figure 4: Model Interpretability Evaluation and Scoring Methodology. (Left) Accuracy of ChemMAS outputs compared to human expert annotations. (Center) Human alignment performance comparison; blue bars indicate LLM-Scores and green bars indicate BLEU-4 scores. (Right) Schematic representation of the LLM-Score pipeline and the question-answering based evaluation workflow.
As shown in Table 2, ChemMAS surpasses both specialized chemical models and state-of-the-art LLMs across all reaction types and Top-$k$ settings. It achieves relative Top-1 similarity improvements ranging from 70% to over 90% when compared to domain-specific baselines such as RCR, Reagent Transformer, and MM RCR. Even against top-tier general-purpose LLMs like GPT-5 and Gemini 2.5-Pro, ChemMAS yields consistent relative gains of 15–25% in Top-1 similarity, underscoring its strength in fine-grained mechanistic reasoning.
5.2 Generalization Evaluation on Out-of-Distribution Data
To rigorously evaluate the generalization capability of our framework and assess chemical reasoning in out-of-distribution (OOD) scenarios, we conducted additional experiments on ChemCoTBench, a standardized benchmark. The primary objective of this experiment is to verify that ChemMAS does not merely rely on retrieving near-duplicate samples from the knowledge base, but truly possesses the ability to perform robust reasoning on novel reaction types.
Specifically, in the challenging Top-1 setting, as shown in Table 1, ChemMAS achieves a significant accuracy advantage. For catalyst prediction, ChemMAS attains an accuracy of 62.1%, surpassing the second-best model (Gemini 2.5-Pro) by a margin of 16.5%. Similarly, for solvent and reagent prediction, ChemMAS outperforms the strongest competitor (GPT5) by 13.7% and 11.1%, respectively. These substantial performance gains on OOD data demonstrate that our ChemMAS framework effectively generalizes beyond the training distribution, exhibiting fine-grained mechanistic reasoning rather than relying solely on memory-based retrieval.
5.3 Evaluation of Model Interpretability
To ensure the interpretability of ChemMAS, we conducted a two-level evaluation focusing on mechanistic grounding and reasoning quality. First, we validated the intermediate outputs of the General Chemist against human ground truth (Figure 4, Left). The agent demonstrates high reliability, achieving accuracies of $95.8\%$ (MainFG), $90.2\%$ (By-product), and $92.5\%$ (Reaction Type), consistently surpassing the $90\%$ threshold. This high alignment confirms that the system builds its downstream reasoning on a correct and verifiable mechanistic understanding.
Building on this foundation, we assessed the quality of the generated reasoning trajectories using a dual-metric framework comprising BLEU-4 and a semantic LLM-Score (Figure 4, Right). The LLM-Score employs an “LLM-as-a-Judge” mechanism to verify if the generated rationale logically supports expert-derived QA pairs. As shown in Figure 4 (Center), ChemMAS significantly outperforms general-purpose LLMs, achieving a superior LLM-Score of $92.8$ compared to the $62.5\text{--}77.2$ range of baselines like DeepSeek-R1 and GPT-5. This substantial gap, alongside a BLEU-4 score of $0.26$, demonstrates that ChemMAS generates scientifically sound explanations rather than merely plausible text.
Table 2: Main results on the private dataset. We report the Top-$k$ similarity (%) across five reaction condition types: catalyst, solvent1, solvent2, reagent1, and reagent2. Results are evaluated at $k\in\{1,5,10\}$. The best and second-best results are bolded and underlined. Green values in parentheses indicate relative improvements over the second-best results.
Top-$k$ Similarity (%)
Model Catalyst Solvent1 Solvent2 Reagent1 Reagent2
@@ 5.3 Evaluation of Model Interpretability
Top-$k$ Similarity (%)
Method Catalyst Solvent 1 Solvent 2 Reagent 1 Reagent 2
@@
Table 4: Ablation study on the SFT, RL, and specific components of the hierarchical reward function, including Acc and $r_{M}$. The best and second-best results are bolded and underlined.
Top-$k$ Similarity (%)
Training Framework Catalyst Solvent 1 Solvent 2 Reagent 1 Reagent 2
@@
Figure 5: Multi-agent ablation: Top-1 similarity improvements across Catalyst, Solvent1/2, and Reagent1/2 when adding specialized agents on top of $\mathcal{A}_{Gen}$$\mathcal{A}_{Full}$.
@@ 5.4 Additional Quantitative Analysis
We conducted an ablation study to analyze the contribution of different components in ChemMAS. The ablation settings are as follows: (1) w/o Main FG, w/o By-Product, and w/o Reaction Type denote removing the corresponding elements from the Memory module; (2) w/o Multi-Agent Debate replaces multi-agent collaboration with a single-agent reasoning process, thereby eliminating conversational exchanges; (3) w/o Multi-Step Reasoning removes the iterative evidence-based reasoning chain within each agent, such that agents can only rely on prior knowledge and inter-agent debate without tool invocation; (4) w/o Candidate Pairing discards the pairwise elimination mechanism for candidate conditions, instead applying a global scoring and ranking procedure to directly select the top-50 candidates. As illustrated in Table 3, removing key components leads to substantial performance drops, underscoring their critical role in ChemMAS. Specifically, removing Main FG from the Memory module results in a significant decrease in performance, with an average drop of +8.4% across all reaction conditions, highlighting the crucial role of functional group extraction and analysis in reaction condition prediction. Similarly, removing Multi-Step Reasoning causes an average similarity decrease of 12.3%, underscoring the importance of evidence-based multi-round reasoning.
To evaluate our framework, we ablated SFT, RL, and specific hierarchical reward components. As shown in Table 4, removing SFT or RL significantly degrades Top-k Similarity across all conditions. Notably, excluding SFT causes a larger drop than removing RL, underscoring the importance of SFT for initialization. We further investigated the reward terms in Eq. (10) by removing task accuracy (Acc) and the multi-tool bonus ($r_{M}$). Results show that ablating $r_{M}$ impairs performance, validating the explicit reward for combined tool usage. Similarly, excluding Acc degrades results, confirming that prioritizing correctness is essential. These findings validate our two-stage framework and hierarchical reward design, where all components play complementary roles.
@@ 6 Conclusion
We introduce ChemMAS, a multi-agent system reframing reaction condition recommendation as evidence-based reasoning grounded in domain knowledge, mechanistic constraints, and interpretable evidence. Unlike prediction-only baselines, ChemMAS explains why conditions are appropriate, enhancing trust and utility. Empirically, it achieves up to 30% Top-1 similarity gains over specialized models and outperforms general LLMs. These results validate the transition from black-box predictions to auditable decision-making. Future work will extend this framework to broader domains such as materials design and bioinformatics.
@@
    • Appendix C.1: Algorithm of ChemMAS Framework and Multi-Agent Debate
@@
  • Appendix D outlines the experimental settings, including:
    • Appendix D.1: Training Pipeline for Agents
    • Appendix D.2: Evaluation Setting Details (Candidate Ranking)
  • Appendix E details the evaluation protocol and metrics:
    • Appendix E.1: SMILES Canonicalization and Validity
    • Appendix E.2: Tanimoto Similarity
    • Appendix E.3: Molecular Fingerprint Types
    • Appendix E.4: Aggregate Evaluation Metrics and Top-k Similarity
  • Appendix F introduces the dataset details:
    • Appendix F: Public Dataset (ChemCoTBench RCR subset)
    • Appendix F: Private Dataset curation and statistics
  • Appendix G presents the prompt templates for different agents.
  • Appendix H presents additional experimental results and discussions:
    • Appendix H.1: Additional Quantitative Results (Top-5 and Top-10 Analysis)
    • Appendix H.2: Result Visualization and qualitative analysis
@@ B.1 Reaction Condition Prediction
Predicting reaction conditions from reactants and products is a long-standing challenge in computer-aided synthesis. Early large-scale efforts such as (Gao et al., 2018) used feedforward neural networks trained on millions of Reaxys records to jointly predict catalysts, solvents, reagents, and temperatures, achieving promising top-k accuracies despite sparsity and label imbalance. Focusing on cross-coupling families, (Maser et al., 2021) formulated the task as multi-label ranking, developing role-specific encoders and leveraging graph-based features to yield accurate, context-aware predictions. To improve generalization and interpretability, (Wang et al., 2023) released benchmark datasets and proposed Parrot, a Transformer model augmented with unsupervised reaction center learning. Parrot achieved significant gains in condition similarity and temperature estimation while offering interpretable attention maps localized to reactive substructures. Separately, (Andronov et al., 2023) addressed data quality limitations by training a Molecular Transformer to impute missing reagents in USPTO reactions. Their system not only improved reagent recall but also enhanced downstream product prediction models.
Retrieval-augmented methods incorporate external knowledge to improve robustness. TextReact (Qian et al., 2023) pairs structure-based encoders with retrieved literature snippets to inform condition prediction and retrosynthesis. By integrating textual context into training, it significantly outperforms molecule-only baselines. In peptide catalysis design, (Edwards et al., 2022) proposed a semi-automated ML framework for selecting universal catalyst libraries and discovered novel, high-selectivity peptides via efficient search in a large tripeptide space. At the interface of language and chemistry, (Edwards et al., 2022) introduced MolT5, a pre-trained encoder-decoder model that translates between molecules and natural language. It supports molecule-to-caption generation and chemically constrained text-to-molecule synthesis, offering a foundation for LLM-based explainability. More recently, (Zhang et al., 2024b) proposed a text-augmented multimodal LLM framework for reaction condition recommendation. Their method jointly encodes SMILES, molecular graphs, and relevant text to achieve state-of-the-art similarity across open benchmarks and improve generalization under low-data or OOD settings. Despite these advances, current methods primarily focus on recommending what the potential reaction conditions are, but fail to provide explanatory why-level evidence for why such conditions are important or mechanistically justified.